Vulcan Company’s contribution format income statement for June is as follows:

Vulcan Company
Income Statement
For the Month Ended June 30
Sales $ 900,000
Variable expenses 400,000
Contribution margin 500,000
Fixed expenses 485,000
Net operating income $ 15,000

Management is disappointed with the company’s performance and is wondering what can be done to improve profits. By examining sales and cost records, you have determined the following:

  1. The company is divided into two sales territories—Northern and Southern. The Northern Territory recorded $400,000 in sales and $160,000 in variable expenses during June; the remaining sales and variable expenses were recorded in the Southern Territory. Fixed expenses of $184,000 and $150,000 are traceable to the Northern and Southern Territories, respectively. The rest of the fixed expenses are common to the two territories.

  2. The company is the exclusive distributor for two products—Paks and Tibs. Sales of Paks and Tibs totaled $290,000 and $110,000, respectively, in the Northern territory during June. Variable expenses are 29% of the selling price for Paks and 69% for Tibs. Cost records show that $145,000 of the Northern Territory’s fixed expenses are traceable to Paks and $26,400 to Tibs, with the remainder common to the two products.

Required:

1-a. Prepare contribution format segmented income statements for the total company broken down between sales territories.

1-b. Prepare contribution format segmented income statements for the Northern Territory broken down by product line.

  • Req 1A

Prepare contribution format segmented income statements for the total company broken down between sales territories. (Round the percentage answers to one decimal place (i.e .1234 should be entered as 12.3).)

Sales Territory
Total Company Northern Southern
Amount % Amount % Amount %
  • Req 1B

Prepare contribution format segmented income statements for the Northern Territory broken down by product line. (Round the percentage answers to one decimal place (i.e .1234 should be entered as 12.3).)

Product Line
Northern Territory Paks Tibs
Amount % Amount % Amount %
Answer #1

1a) Segmented income Statement for the Total Company Sales Teritory Total Company Amount Northren South % Amount | % |Amount | % Sales Variabel Expenses Contribution Margin Traceable Fixed Expenses Market Segmented Margin Common Fixed Cost Net operating Income 900000 100.0% 400000 100.0% 500000 100.0% 400000 44.4% 160000 40.0% 240000 48.0% 500000 55.6% 240000 - 400% 150000 . 30,0%) 60.0% 260000 52.0% 334000 37.1% 184000 46.0% 150000 30.0% 166000 151000 16.8% 18.4% 56000 14.0% 110000 22.0% 15000 1.7901b) Segmented income Statement for the Northren Terrotory Product Lines Northern Territorv Paķs Tibs Amount % Amount | % |Amount | % Sale Variabel Expenses Contribution Margin Traceable Fixed Expenses Market Segmented Margin common Fixed Cost Net operating Income 400000 100.0% 290000 100.0% 110000 100.0% 160000 40.0% 84100 29.0% 75900 69.0% 240000 60.0% 205900 71.0% 34100 31.0% 171400 42.9% 145000 50.0% 26400 24.0% 68600 17.2% 60900 21.0% 7700 7.0% 12600 3.2% 56000 14.0%

Total Sales = Northen sales + Sothern Sales

900,000 = 400,000 +500,000

Sothern Varaible cost = Total Variable cost - Northen Variable cost = 500000-160000= 340000

Traceable Fixed cost: 334,000 , Northen 184000 Sothern 150,000

Common Fixed Cost =485,000 -334,000 =151,000

Variable cost of Paks = 290,000 * 0.29 = 84,100

Varaible cost of Tibs = 110,000 *0.69 =75,900